OpenStudy (anonymous):

hhhheeeeeeeellllllllllp..................... Find the equation of the line below. Arrange your answer in the form y = mx + b, where b is the constant. (Note: in this answer make sure your x has a coefficient) (2,1)(-2,-1)

OpenStudy (anonymous):

slope of line ( m ) = y2 - y1/x2 - x1 , tell me what u get

OpenStudy (anonymous):

m=-2/-4

OpenStudy (anonymous):

yep , that's 1/2 (simplified) now to find b , use the equation y = mx + b choose any point and plug in the values , now that u have m

OpenStudy (anonymous):

1=1/2*2+b

OpenStudy (anonymous):

b=0 and i need a fraction....

OpenStudy (anonymous):

y-y1=m(x-x1) y-1=1/2(x-2) 2y-2=x-2 2y=x y=1/2x+0 This line passes through (0,0) so constant b=0

OpenStudy (anonymous):

but my worksheet says it need to be a fraction...

OpenStudy (anonymous):

check the points given.

OpenStudy (anonymous):

how do u mean?????

OpenStudy (anonymous):

I mean to say they may be different e.g.,(2,-1) or (-2,1) etc.

OpenStudy (anonymous):

is the answer for this problem 0?????

OpenStudy (anonymous):

Answer as given in the question is y=1/2 x+0

OpenStudy (anonymous):

ok, i get it,,, thank u soooo much....ps. what does coefficient mean..

OpenStudy (anonymous):

In y=1/2 x+0 co efficient of x=1/2 coefficient of y=1 ax+by+c=0 coefficient of x=a coefficient of y=b

OpenStudy (anonymous):

thank u soo much.....